OpenStudy (anonymous):

**WILL MEDAL** Larry's Landscaping charges $235 for spring cleanups and $30 for weekly lawn maintenance. Joe's Landscaping charges $185 for spring cleanups and $35 for weekly lawn maintenance. The system that models this situation is given, where c is the cost of lawn maintenance and w is the number of weeks. c = 235 + 30w c = 185 + 35w The solution to the system is (10, 535). Which interpretation correctly describes the solution to the system of equations?

OpenStudy (anonymous):

A. Larry’s Landscaping will charge more money on the tenth week by charging $535. B. Joe’s Landscaping will charge more money on the tenth week by charging $535. C. The cost for lawn maintenance is the same, $535, for both landscaping companies after 10 wk. D. A customer can have his or her lawn maintained for a maximum of 10 wk. The customer will pay a total of $535.

OpenStudy (tylerd):

so the solution is (10,535) we have equations in the form c=235+30w and c=235+35w this is a function c(w) w is our "x" so the point (10,535) w=10 and c=535

OpenStudy (anonymous):

So D?

OpenStudy (tylerd):

the 2 lines apparently intersect at (10,535) so the cost is the same at 10 weeks

OpenStudy (anonymous):

Oh ok C then

OpenStudy (tylerd):

ya

OpenStudy (anonymous):

Thx